Solve the system of equations.
1.25x – 2y = 1.3125
x + 2y = –0.75

A) {(1.75, 1.5)}
B) {(0.75, 0)}
C) {(0.25, –0.5)}
D) {(1.75, –0.5)}

Answer by stanbon(75887)

Solve the system of equations.
1.25x – 2y = 1.3125
x + 2y = –0.75
------------------------
Add the two equations to get:
2.25x = 0.5625
x = 0.25
------------------
Substitute to solve for "y":
0.25 + 2y = -0.75
2y = -1
y = -1/2
==============
Cheers,
Stan H.